# Content ARCs: Decentralized Content Rights in the Age of Generative AI

### v2

@@ 3.1 Authenticity
Metadata schemes provide a way to embed a unique identifier (GUID) in an asset, for example: a Manifest ID in C2PA; or EIDR, for IDs in audiovisual content. Critics of metadata schemes, point to their fragility due to metadata stripping on content platforms (c.f. section 2), motivating the need to ground the identifier in a durable identifier derived from content itself (‘content ID’), such as a watermark or fingerprint. In contrast to the open nature of metadata standards, content ID schemes are often proprietary. This presents a challenge to ARCs in enabling a low-friction, interoperable solution for robust decentralized rights management. Recently, the International Standard Content Code (ISCC)[14] became an international standard (ISO 24138), describing within a standard method to encode image fingerprints. However, the method adopted is a decades-old perceptual hash (pHash [21]) that exhibits a high false positive matching rate [15] at global scale (i.e. over trillions of assets). This is due to the simplistic encoding method based on frequency analysis (DCT) which cannot discriminate smaller differences in images, and its small (64 bit) key space. The cryptographic ‘birthday attack’ estimates the probability of a collision in $n$ bits as $P(n)\approx 1-e^{-\frac{n^{2}}{2^{65}}}$ i.e. a 50% chance at only 5 billion assets for 64 bits. Other open technologies based on DCT are PhotoDNA [22] (144 bits, mainly used for CSAM identification) and PDQ [23] (256 bits) however these still produce false matches under small manipulations. Modern, performant methods for content ID rely upon a trained neural network (AI) to derive content ID e.g. [24, 15, 13]. Herein lies a further paradox. Open standards are the only practical solution to adoption and interoperability. However the pace of standardization is at odds with fast moving emerging technologies like content ID; and functionality defined by a trained AI model is difficult to standardize. Similar arguments may be made for watermarking, where older methods based on frequency domain processing are being superseded by robust AI approaches [11]. This is leading toward de facto standardization practices where open sourced content ID models are being commercially adopted (e.g. TrustMark [11]), creating silos of interoperable solutions.
@@ 3.3 Compensation
Attribution-Driven Payouts. The training datasets responsible for creating AI models can be described via provenance metadata attached to models, raising the opportunity to pay royalties to contributors for use of the model e.g. when synthetic assets are generated [18]. However, for Generative AI the scale of such datasets runs into the billions, presenting a challenge to fractional compensation. Data attribution can help identify the subset of training data most influential in a given synthetic asset, to target payout. Several techniques have measured attribution as visual correlation between output and training data [28, 29]. Yet correlation does not equal causation. Causative methods have been proposed, such as invisible watermarking in ProMark[30] to measure which mixture of watermarks exists in the output, as well as Shapley value-based analytics [31], however neither scales to billions of data items. Scaling causative attribution techniques to the billion-scale datasets used for training GenAI models remains an open research challenge in the development of royalty mechanisms for GenAI.

### v3

@@ 3.1 Authenticity
Metadata schemes provide a way to embed a unique identifier (GUID) in an asset, for example: a Manifest ID in C2PA; or EIDR, for IDs in audiovisual content. Critics of metadata schemes, point to their fragility due to metadata stripping on content platforms (c.f. section 2), motivating the need to ground the identifier in a durable identifier derived from content itself (‘content ID’), such as a watermark or fingerprint. In contrast to the open nature of metadata standards, content ID schemes are often proprietary. This presents a challenge to ARCs in enabling a low-friction, interoperable solution for robust decentralized rights management. Recently, the International Standard Content Code (ISCC)[14] became an international standard (ISO 24138:2024), describing within a standard method to encode ‘image codes’ using the perceptual hash (pHash [21]); a frequency based (DCT) algorithm. This may be modified from its usual 64 bits up to 256 bits by concatenating the result computed over a $2\times 2$ spatial grid [22]. The cryptographic ‘birthday attack’ estimates the probability of a collision in $n$ bits as $P(n)\approx 1-e^{-\frac{n^{2}}{2^{65}}}$ i.e. a 50% chance at only 5 billion assets for 64 bits, making such a facility important at global scale (i.e. over trillions of assets). Other open technologies based on DCT are PhotoDNA [23] (144 bits, mainly used for CSAM identification) and PDQ [24] (256 bits). In general DCT hashes exhibit a high false positive matching rate [15], producing false matches under small manipulations. Contemporary methods for content ID rely upon a trained neural network (AI) to derive a descriptor e.g. [25, 15, 13]. Herein lies a further paradox. Open standards are the only practical solution to adoption and interoperability. However the pace of standardization is at odds with fast moving technologies like content ID; and functionality defined by a trained AI model is difficult to standardize. Similar arguments may be made for watermarking, where methods based on frequency domain processing are being superseded by AI approaches [11]. This is leading toward de facto standardization where open sourced content ID models are being commercially adopted (e.g. TrustMark [11]), creating silos of interoperable solutions. Some approaches combine both watermarking and perceptual hashing for additional durability [18]. Whilst the above focused on images, the same holds for other modalities.
@@ 3.3 Compensation
Attribution-Driven Payouts. The training datasets responsible for creating AI models can be described via provenance metadata attached to models, raising the opportunity to pay royalties to contributors for use of the model e.g. when synthetic assets are generated [18]. However, for Generative AI the scale of such datasets runs into the billions, presenting a challenge to fractional compensation. Data attribution can help identify the subset of training data most influential in a given synthetic asset, to target payout. Several techniques have measured attribution as visual correlation between output and training data [29, 30]. Yet correlation does not equal causation. Causative methods have been proposed, such as invisible watermarking in ProMark[31] to measure which mixture of watermarks exists in the output, as well as Shapley value-based analytics [32], however neither scales to billions of data items. Scaling causative attribution to billion-scale datasets for GenAI models remains an open research challenge in the development of royalty mechanisms for GenAI.